OpenStudy (firejay5):

The direction of the 60-pound force is north, and the direction is east for the 100-pound force. Describe the direction of the net force.

OpenStudy (firejay5):



OpenStudy (bioepic):

srry gtg @PRAETORIAN.10

OpenStudy (praetorian.10):

...............physics

OpenStudy (praetorian.10):

ummmm

OpenStudy (praetorian.10):

isnt it just 160 North East

OpenStudy (firejay5):

Yea can you help the teacher said the answer was 59 degrees Northeast

OpenStudy (praetorian.10):

yeah bro i am like the last person you wanna ask about math/physics

OpenStudy (praetorian.10):

you may as well ask a stick it probs knows more than me

OpenStudy (praetorian.10):

but i would say the answer to get 59 would be 100 - 100 plus 60 and then take 1/100 = 1 so 59

OpenStudy (firejay5):

I thought I would try, if you need help with english/literature/science; I'll surely help you

OpenStudy (praetorian.10):

so yeah i dunno if that helps buddy

OpenStudy (praetorian.10):

but it does equal 59 in the end

OpenStudy (firejay5):

How does it

OpenStudy (praetorian.10):

but i would say the answer to get 59 would be 100 - 100 plus 60 and then take 1/100 = 1 so 59

OpenStudy (praetorian.10):

(i wrote it above)

OpenStudy (firejay5):

like expand on it

OpenStudy (praetorian.10):

lol i dunno mate

OpenStudy (firejay5):

I get 100 - 100 = 0 + 60 = 60 part

OpenStudy (praetorian.10):

then you just take 1/100 because there are two 100's in the equation

OpenStudy (praetorian.10):

i guess

OpenStudy (praetorian.10):

i dunno

OpenStudy (praetorian.10):

i suck

OpenStudy (praetorian.10):

i bet some maths people know this

OpenStudy (praetorian.10):

im a history guy

OpenStudy (firejay5):

@iambatman @Kainui

OpenStudy (praetorian.10):

sorry i couldnt be much help

OpenStudy (praetorian.10):

ask me about ancient roman guys anyday

OpenStudy (perl):

the direction is tan theta = 60/100 arctan (60/100) = theta

OpenStudy (perl):

theta = 30.96 degrees

OpenStudy (perl):



OpenStudy (praetorian.10):

yep.....like i said xD

OpenStudy (firejay5):

@PRAETORIAN.10 no you didn't say that, because like you said, "you are the last person to ask"

OpenStudy (praetorian.10):

yeah it was a joke hence the laughing face......